OpenStudy (safari321):

You deposit $200 in an account that earns 5% simple interest. How long will it be before the total amount is $400?

OpenStudy (mathmate):

approximately 14.4 years using the rule of 72 (72/5=14.4). To find the exact number of years, use the compound interest formula FV=PV*(1+r)^t from which t=log(FV/PV)/log(1+r) =14.21 years.